OpenStudy (anonymous):

Simplify 6/sqrt 3 +2

OpenStudy (anonymous):

\[\frac{ 6 }{ \sqrt{3}+2 } ?\]

OpenStudy (anonymous):

multiply and divide by the conjugate that is \[\sqrt{3} - 2\]

OpenStudy (anonymous):

and then solve as \[(6\div \sqrt{3}+2) *(\sqrt{3}-2 \div\sqrt{3} - 2)\]

OpenStudy (solomonzelman):

yes, use conjugate. (The purpose of it is to avoid the middle term, because what you will get in the denominator is something similar to " (a+b)(a-b)" which simplifies to "a^2-b^2" - without any middle term. The reason we are trying to avoid the middle term is because we can't have a radical in the denominator, but we will if we get any middle term.) \(\huge\color{blue}{ \frac{ 6 }{ \sqrt{3} +2 } }\) follow me from here. \(\huge\color{blue}{ \frac{ 6 }{ \sqrt{3} +2 } }\) \(\huge\color{red}{ \times \frac{ \sqrt{3}-2 }{ \sqrt{3} -2 } }\) \(\huge\color{blue}{ \frac{ 6 ~\color{red} {(~\sqrt{3} -2~) } }{ (\sqrt{3} +2)~\color{red} {(~\sqrt{3} -2~) } } }\) \(\huge\color{blue}{ \frac{6\sqrt{3}-12}{\sqrt{3}^2-\sqrt{2}^2} }\) for the bottom I used (a+b)(a-b)=a^2-b^2 and for the top just expended. Now, lets finish simplifying... \(\huge\color{blue}{ \frac{6\sqrt{3}-12}{9-4}}\) \(\huge\color{blue}{ \frac{6\sqrt{3}-12}{5}}\) THAT'S THE SIMPLEST FORM \(\color{green}{ :)}\)